# Illuminating Darkness: Enhancing Real-world Low-light Scenes with Smartphone Images

### v1

@@ Abstract
Digital cameras often struggle to produce plausible images in low-light conditions. Improving these single-shot images remains challenging due to a lack of diverse real-world pair data samples. To address this limitation, we propose a large-scale high-resolution (i.e., beyond 4k) pair Single-Shot Low-Light Enhancement (SLLIE) dataset. Our dataset comprises 6,425 unique focus-aligned image pairs captured with smartphone sensors in dynamic settings under challenging lighting conditions (0.1–200 lux), covering various indoor and outdoor scenes with varying noise and intensity. We extracted and refined around 180,000 non-overlapping patches from 6,025 collected scenes for training while reserving 400 pairs for benchmarking. In addition to that, we collected 2,117 low-light scenes from different sources for extensive real-world aesthetic evaluation. To our knowledge, this is the largest real-world dataset available for SLLIE research. We also propose learning luminance-chrominance (LC) attributes separately through a tuning fork-shaped transformer model to enhance real-world low-light images, addressing challenges like denoising and over-enhancement in complex scenes. We also propose an LC cross-attention block for feature fusion, an LC refinement block for enhanced reconstruction, and LC-guided supervision to ensure perceptually coherent enhancements. We demonstrated our method’s effectiveness across various hardware and scenarios, proving its practicality in real-world applications. Code and dataset available at https://github.com/sharif-apu/LSD-TFFormer.
@@ 1 Introduction
Low-light photography has gained significant popularity due to its aesthetic appeal and widespread availability of compact handheld devices. However, digital cameras usually suffer quantum inefficiency and increase noise levels in low-light conditions, reducing image quality. These challenges have fueled an unprecedented demand for effective learning-based SLLIE techniques. Beyond aesthetic enhancement, SLLIE has practical applications across various domains, including surveillance, autonomous systems, and general vision tasks such as image matching and object detection (OD) [62, 59, 37].
Learning-based SLLIE remains a challenging problem due to the scarcity of high-quality training datasets and the difficulty of benchmarking solutions across diverse real-world conditions [17]. Notably, the existing SLLIE datasets [49, 55, 13, 27] are often constrained by controlled environments, limited dynamic settings, and semi-synthetic low-light conditions created using fixed ISO values (e.g., ISO 50, 100 in LSRW [13]). In contrast, real-world low-light imaging involves higher ISO values and faster shutter speeds, leading to significant noise, blur, and color distortions [37, 64]. Current datasets fail to capture these complexities, hindering deep models from effectively generalizing to real-world low-light conditions. As a result, deep models tend to amplify noise, introduce unnatural color shifts, and generate artifacts in challenging scenes, as shown in Fig. 1.
Figure 1: Real-world complex scene enhancement using SLLIE datasets. From left to right: Input, LOL-V1 [49], LOL-V2 [55], LSRW [13], and NTIRE challenge on low-light enhancement (NCLLIE) [27], and the proposed LSD. We trained the proposed TFFormer separately on each dataset to evaluate their performance in handling real-world complex scenarios. The proposed LSD helps the model better understand the scene and enhances low-light inputs, avoiding the noise amplification and visual artifacts seen in existing methods.
Despite data limitations, recent research has explored different SLLIE [17] approaches. The end-to-end learning-based methods [2, 29, 43, 39, 26, 30, 22, 53, 11, 9, 15] aim to directly map low-light images to well-lit counterparts, often integrating Convolutional Neural Networks (CNNs). Several recent methods leverage transformers [52, 7, 41] to enhance the receptive field of vanilla convolutions for learning spatial information, inspired by advances in image restoration [57, 48, 58]. Contrastly, Retinex-based methods [19, 4, 49, 56, 50, 61, 46, 62, 47, 40, 60, 63] incorporate Retinex theory into CNN architectures. Retinex-CNN methods typically decompose RGB images, denoise reflectance, and adjusts illumination leveraging separate networks. A few recent methods [4, 33, 35] also explored hybrid models combining CNNs, transformers, and Retinex theory to improve visual reconstruction while performing efficient training. Regrettably, these methods can produce desaturation and brightness inconsistencies similar to Retinex-CNN approaches[35]. In sum, SLLIE is hindered by limited datasets and model shortcomings, including noise amplification, artifacts, and color distortions in real-world conditions.
In this paper, we tackle the real-world SLLIE by introducing a large-scale dataset and a novel deep network. Our low-light smartphone dataset (LSD) comprises 6,425 high-resolution, focus-aligned low-light images with corresponding ground truth. We captured all scenes leveraging dynamic settings and lighting conditions ranging from 0.1 to 200 lux. To ensure diversity, we collected indoor and outdoor image pairs featuring noisy and denoised low-light inputs. We extracted 179,694 patches from 6,025 scenes for training and used the remaining 400 scenes for testing. Additionally, for generalizing SLLIE in real-world scenarios, we collected an unpair benchmarking dataset (denoted as LSD-U in later sections) of over 2,100 unique scenes sourced from smartphones, DSLRs, video frames, social media, etc.
We also introduce TFFormer, a tuning fork (TF)-shaped Transformer-CNN model [3, 41, 48] to address learning real-world SLLIE. Taking inspiration from human vision [16, 31], our proposed network learns luminance and chrominance separately instead of using Retinex-based reflection-illumination decomposition. Unlike multi-stage Retinex networks, our streamlined single-network architecture offers a more straightforward design while addressing over-enhancing (i.e., brightness, desaturation, etc.) denoising. It also improves independent noise reduction in luminance and chrominance through the LC attribute. To leverage the LC attributes precisely, we introduce a Luminance-Chrominance Cross-Attention Block (LCCAB) to correlate encoded LC attributes for LC fusion and global consistency. Additionally, we introduce a Luminance-Chrominance Guided Refinement Block (LCGRB) to enhance reconstructions further. To ensure perceptually coherent enhancements, we propose a Luminance-Chrominance Perceptual Loss (LCPL) to guide our network toward generating visually plausible results. The proposed TFFormer outperforms the existing methods in real-world complex SLLIE. We generalize our method in diverse scenarios. Our contributions are as follows:
  1. 1.
    We collected the largest SLLIE dataset using smartphones. The proposed LSD includes 6,425 focus-aligned high-resolution low-light image pairs and an unpaired benchmark of 2,117 unique scenes from diverse sources.
  2. 2.
    We introduced a TFFormer to outperform existing methods in real-world SLLIE. Our proposed network leverages LC priors along with LCCAB, LCGRB, and LCPL to perceive visually plausible images.
  3. 3.
    We generalized and illustrated the practicability of our method across numerous hardware and use cases. Our evaluation included multiple user studies, perceptual comparisons, quantitative comparisons, and impacts on vision tasks.
2 Low-light Smartphone Dataset (LSD)
2.1 SLLIE Datasets
Four real-world paired datasets have been widely used in existing real-world SLLIE: Low-light (LOL) V1 [49], LOL-V2 [55], LSRW [13], and NCLLIE [27]. Table 1 summarizes the characteristics and sample counts of SLLIE datasets. Notably, the existing datasets were collected in controlled environments and used only a limited range of image sensors. Consequently, they lack diversity regarding light variation and noise, as illustrated in Fig. 2. This scarcity of unique scenes and varied benchmarking samples has significantly constrained progress in SLLIE. Our proposed LSD overcomes these limitations by providing a high-quality dataset specifically designed to tackle the challenges associated with real-world SLLIE. As shown in Fig. 2, our LSD incorporates a broader distribution of light intensity and noise characteristics in the input images and their corresponding ground truth, offering a more diverse and realistic representation of real-world low-light scenarios compared to existing datasets.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ 2.1 SLLIE Datasets
Figure 2: Noise and light intensity distribution comparison between LSD and existing SLLIE datasets. LSD dataset demonstrates higher noise diversity and light intensity variation than existing SLLIE datasets, enhancing SLLIE methods’ generalization.
@@ 2.2 Scene acquisition
Figure 3: Distribution and sample images of proposed LSD. (a) DLI and NLI distribution, (b) DLI and NLI sample image pairs.
Scene Setup. Collecting real-world paired data for SLLIE is significantly challenging, especially in obtaining perfectly aligned image pairs without motion artifacts or spatial disparity. Despite the challenges, we captured over 8,000 scenes featuring low-light and well-lit reference images. The data was collected in uncontrolled outdoor environments—from sunset to sunrise—using 15 camera setups across six modern smartphones. These smartphones offered diverse camera configurations and allowed for precise control of settings, making them an ideal choice for large-scale data collection. We incorporated a three-frame strategy to ensure spatial alignment, sandwiching each low-light image between two reference images. A proprietary camera app, developed with the Android Camera API [12], controlled sensitivity, shutter speed, and denoising while other settings remained auto. In each scene, reference images were denoised using the smartphone’s ISP. At the same time, low-light inputs were collected in two versions: (1) Denoised Low-Light Inputs (DLI) and (2) Noisy Low-Light Inputs (NLI), introducing noise diversity. We retained only scenes with at least one perfectly aligned low-light input and reference image, enhancing visibility with histogram matching [36, 38]. Fig. 3(b) shows sample images from NLI and DLI.
@@
Capture Calibration. We employed a dynamic calibration strategy to capture low-light inputs and their corresponding reference images for introducing dataset’s robustness and diversity. For reference images, settings were calibrated to a low ISO range (i.e., 50–400) to minimize sensor noise. To achieve bright, naturally lit reference images, we used long exposure times, ranging from 1,000 to 15,000 ms, depending on the scene conditions. For low-light inputs, we utilized two distinct capture settings:
  1. 1.
    ISP-driven settings: We took a dummy image in the stock camera app using auto mode before capturing each low-light scene. We then analyzed the sensitivity and shutter speed used in the dummy capture and applied these settings to the low-light input image to replicate camera characteristics.
  2. 2.
    Random-sensitivity settings: ISO and shutter speed configurations are hardware-dependent[20]. However, we aimed to generalize SLLIE across various scenes and devices. Thus, in approximately 60% of scenes, we randomly adjusted ISO and shutter speed to introduce noise diversity while maintaining a low-light scene appearance similar to the dummy capture.
@@
LSD distribution: From over 8,000 collected scenes, we selected 6,425 aligned-focus scenes to construct the LSD dataset, comprising high-resolution image pairs ranging from $3820\times 2160$ to $4080\times 3060$. The dataset includes 2,893 indoor and 3,532 outdoor scenes captured under diverse lighting conditions from 0.1 to 200 lux, as illustrated in Fig. 3(a). We categorized lighting into extreme low-light (0.1–50 lux) and low-light (50–200 lux) conditions. Of these, 6,025 pairs were used for training, while 400 complete scenes were selected for benchmarking, including DLI and NLI samples across various lighting conditions in both indoor and outdoor settings. Further details are provided in the supplementary material
@@
Figure 4: Distribution comparison between filtered and raw training images. (a) intensity filtering. (b) sharpness improvement (blur mitigation).
@@ 2.3 Refining Training Pairs
Despite dynamic settings in uncontrolled environments, reference images may exhibit minor imaging limitations, such as defocus blur, absolute dark regions, and overexposed light sources [20, 24] in localized areas. These issues can occasionally cause deep models to produce blurry regions and inconsistent exposure in extreme cases (see Sec. 4.4). To mitigate this, we cropped non-overlapping patches of size $512\times 512$ from the reference and low-light images, yielding 208,853 unique training pairs. We eliminated approximately 14% of imperfect pairs by observing the characteristics of the reference images, resulting in 179,695 refined pair patches for learning effective SLLIE. Specifically, the pairs whose reference images have an average brightness below a threshold were discarded. The average brightness $I_{\text{avg}}$ of a reference image is calculated as:
$I_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}I_{i}\quad\text{where}\quad I_{\text{avg}}<T=10$ (1)
where $N$ is the total number of pixels in the image and $I_{i}$ is the intensity of the $i$-th pixel.
Additionally, we developed a VGG-based [25] scene classifier to identify and exclude imperfect pairs. The confidence score $C$ for each reference image, computed as $C=f(\mathbf{I_{G}})$ (where $\mathbf{I_{G}}$ represents the reference image), was used to filter out pairs with $C<0.90$. Since reference images guide deep models in generating plausible outputs, we prioritized their quality during filtering. We also visually inspected the discarded patches to ensure no valuable pairs were lost. Fig. 4 compares the distributions of the raw and filtered patches, highlighting the improvements in intensity and sharpness in the refined training set.
@@ 2.4 LSD-U: Unpair Benchmarking Dataset
SLLIE is inherently a matter of aesthetic evaluation, requiring methods to generalize effectively across real-world scenarios. Therefore, the evaluation should involve diverse scenes with a broad distribution rather than relying on datasets with specific biases. To support a comprehensive subjective evaluation in SLLIE, we introduced the LSD-U, including 2,117 unique unpaired test cases. Our benchmark features data from DSLR cameras, smartphones, low-light video frames, and social media images sourced from various users with different photographic styles and setups compared to the proposed LSD. The LSD-U distribution introduces varied biases, enabling SLLIE methods to be tested across a wide range of real-world conditions, including enhancing trillions of social media images.
@@ 3 Learning with TFFormer
TFFormer maps a low-light image ($\mathbf{I_{\mathbf{LL}}}$) to an intermediate reconstruction ($\mathbf{I_{rec}}$) and a refined output ($\mathbf{I_{ref}}$) as $\mathrm{T}:\mathbf{I_{\mathbf{LL}}}\to{\mathbf{I_{rec}},\mathbf{I_{ref}}}$. Therefore, we first separated LC and encoded them with a dedicated encoder. The encoded LC features are correlated to capture their dependencies. These encoded-refined features are later decoded to generate an intermediate reconstructed image ($\mathbf{I_{rec}}$). We refined $\mathbf{I_{rec}}$ further to produce $\mathbf{I_{ref}}$ leveraging LCGRB. Fig. 5 illustrates the overview of the proposed TFFormer.

### v2

# Illuminating Darkness: Learning to Enhance Low-light Images In-the-Wild

@@ Abstract
Single-shot low-light image enhancement (SLLIE) remains challenging due to the limited availability of diverse, real-world paired datasets. To bridge this gap, we introduce the Low-Light Smartphone Dataset (LSD), a large-scale, high-resolution (4K+) dataset collected in the wild across a wide range of challenging lighting conditions (0.1–200 lux). LSD contains 6,425 precisely aligned low and normal-light image pairs, selected from over 8,000 dynamic indoor and outdoor scenes through multi-frame acquisition and expert evaluation. To evaluate generalization and aesthetic quality, we collect 2,117 unpaired low-light images from previously unseen devices. To fully exploit LSD, we propose TFFormer, a hybrid model that encodes luminance and chrominance (LC) separately to reduce color-structure entanglement. We further propose a cross-attention-driven joint decoder for context-aware fusion of LC representations, along with LC refinement and LC-guided supervision to significantly enhance perceptual fidelity and structural consistency. TFFormer achieves state-of-the-art results on LSD (+2.45 dB PSNR) and substantially improves downstream vision tasks, such as low-light object detection (+6.80 mAP on ExDark).
Figure 1: Qualitative comparison of real-world low-light enhancement using our proposed TFFormer trained on different datasets. From left to right: Input image, TFFormer trained on LOL-V1 [42], LOL-V2 [49], LSRW [12], NCLLIE [25], and the proposed LSD dataset. While existing datasets lead to artifacts or noise amplification, LSD enables robust enhancement with preserved structure and color fidelity, demonstrating its superior diversity and realism.
@@ 1 Introduction
Learning-based SLLIE has gained significant momentum in recent times. It can improve the aesthetic appearance and significantly improve the performance of widely used vision tasks [55, 52, 32]. Unfortunately, models trained on existing SLLIE datasets, such as low-light (LOL) -V1 [42], LOL-V2 [49], LSRW [12], and NTIRE challenge low-light enhancement (NCLLIE) [25], struggle to generalize to complex real-world environments. As a result, deep models often produce over-brightened areas with color shifts (Fig. 1, top) and amplified noise (Fig. 1, bottom).
The primary limitation of SLLIE stems from the characteristics of the training datasets. Existing datasets offer a limited number of training and evaluation samples captured in controlled environments [44]. These datasets often rely on synthetic low-light images created with low ISO settings (e.g., ISO 50/100 in LSRW [12]) and artificially dimmed bright-light scenes. Therefore, they fail to represent critical real-world phenomena: sensor-induced noise, spatially varying light intensity, and color distortions that usually arise from the high-ISO amplification and short exposure times utilized to mitigate motion blurs in stochastic lighting conditions[32, 57]. The constrained noise and intensity distributions of these datasets are illustrated in Fig. 2, highlighting their significantly limited dynamic range.
@@
Figure 2: Noise and intensity distribution comparison between LSD and prior SLLIE datasets. (a) light intensity variation. (b) noise variation.
To address these limitations, we introduce LSD, a large-scale real-world dataset comprising 6,425 high-resolution, focus-aligned image pairs captured using 15 different smartphone camera setups. The data spans diverse indoor (2,893) and outdoor (3,532) environments under uncontrolled lighting conditions ranging from 0.1 to 200 lux. Of these, 6,025 scenes are used to extract 179,694 training patches, while 400 complete scenes are reserved for evaluation. To the best of our knowledge, LSD is the largest publicly available real-world SLLIE dataset in terms of full-scene, high-resolution image pairs (see Table 1 for details). Notably, LSD offers significantly greater scene and sensor diversity compared to the previously largest SLLIE dataset, LSRW [12], with 5,600 cropped patch pairs. To evaluate generalization beyond paired training, we introduce LSD-U, a set of 2,117 unpaired low-light images from unseen devices with diverse imaging characteristics. Models trained on LSD demonstrate strong generalization and substantially improve downstream tasks, achieving a +6.80 mAP gain in object detection on the ExDark dataset [26].
@@
While LSD addresses dataset-related limitations, existing SLLIE models also struggle under real-world degradations. RGB-based models often [40, 45, 51, 41, 6, 13, 27, 37, 34, 24, 28, 19, 46, 10, 8, 15] entangle semantic structures with signal-dependent noise, leading to color instability and texture loss. Retinex-inspired methods [42, 55, 40, 3, 50, 54, 39, 35, 53, 56] decouple illumination and reflectance but often misestimate illumination in complex scenes, resulting in halo artifacts. Recent non-RGB methods (e.g., HVI [47], LYT [2]) attempt to separate luminance and chrominance, but they suffer from hue discontinuities and chroma noise due to limited cross-domain interaction. Please see Sec. 4.1 for more detailed comparisons.
We propose TFFormer, a tuning-fork-shaped Transformer-CNN architecture that addresses key limitations in existing SLLIE methods. Unlike RGB-based [41, 6, 13] and Retinex-inspired [3, 50] models, TFFormer encodes luminance and chrominance separately to enable domain-specific feature learning for structure-preserving denoising and color-consistent enhancement. To bridge the inter-domain correlation gaps observed in recent fixed color-space methods [47], we introduce a Luminance–Chrominance Cross-Attention Block (LCCAB) and a joint decoder that enables adaptive fusion exclusively at the decoding stage. This late-stage interaction leverages spatially adaptive fusion, preserving domain-specific features while ensuring a structurally coherent reconstruction. Furthermore, we introduce the LC Guided Refinement Block (LCGRB) and LC Consistency Guidance loss (LCCG) to improve global structural fidelity and perceptual alignment. TFFormer outperforms the prior SOTA by +2.45 dB PSNR on LSD and generalizes well across diverse real-world scenes.
Contributions. Our main contributions are:
  1. 1.
    We present LSD, the largest real-world SLLIE dataset, featuring over 6,400 paired samples and a 2,100-image unpaired benchmark for generalization testing.
  2. 2.
    We propose TFFormer, a novel LC-decoupled hybrid Transformer-CNN architecture featuring LCCAB, LCGRB, and LCCG for perceptually consistent enhancement.
  3. 3.
    We validate our method extensively with user studies, downstream vision task improvement, and generalization across unseen scenes.
Figure 3: Overview of the proposed LSD dataset: (a) collection and refinement pipeline, (b) Illumination distribution across indoor and outdoor scenes, (c) DLI and NLI sample pairs.
2 Low-light Smartphone Dataset (LSD)
We present LSD, the largest SLLIE dataset collected in-the-wild. Fig. 3(a) illustrates the LSD collection pipeline.
@@ 2.1 Scene acquisition
Collection Setup. Capturing real-world aligned image pairs for SLLIE is inherently challenging due to motion artifacts and spatial disparities [44]. We collected more than 8,000 scenes using 15 camera setups. The scenes were captured in uncontrolled environments, spanning from sunset to sunrise. To fully exploit the proposed pipeline (Fig. 3(a)), we developed an Android app with adjustable ISO, shutter speed, and denoising, leaving other settings on auto. To ensure spatial alignment, we adopted a three-frame capture strategy where two high-quality references bracketed each low-light image. To mitigate the effects of transient motion, we waited for stable moments, such as pauses between moving clouds or vehicles, before initiating capture. Additionally, our multi-frame setup allowed us to verify and reject scenes with observable motion inconsistencies. We acquired two input variants: (1) denoise low-light input (DLI): applied denoising to low-light images processed by the smartphone ISP, and (2) noisy low-light input (NLI): raw noisy captures for enhancing diversity.
@@
Capture Calibration. We employed a dynamic calibration strategy to capture low-light inputs and their corresponding reference images. For reference images, settings were calibrated to a low ISO range (i.e., 50–400) to minimize sensor noise. To achieve bright, naturally lit reference images, we used long exposure times, ranging from 1,000 to 15,000 ms, depending on the scene conditions. For low-light inputs, we utilized two distinct capture settings: (1) ISP-guided: ISP-driven settings: We took a dummy image in the stock camera app using auto mode before capturing each low-light scene. We then analyzed the sensitivity and shutter speed used in the dummy capture and applied these settings to the low-light input image to replicate camera characteristics. (2) Random-sensitivity settings: ISO and shutter speed configurations are hardware-dependent[17]. However, we aimed to generalize SLLIE across various scenes and devices. Thus, in approximately 60% of scenes, we randomly adjusted ISO and shutter speed to introduce noise diversity while maintaining a low-light scene appearance similar to the dummy capture.
Align-pair Selection. For each scene, we matched the low-light input with two adjacent reference frames, generating two candidate pairs. We found that extreme low-light conditions significantly degraded the visibility of low-light inputs, making automatic alignment unreliable. To address this, we enhance the visibility of low-light images (for aligning purposes only) using histogram matching [31, 33], followed by an expert manual review to identify well-aligned pairs. Based on expert assessment, we discarded approximately 1,500 scenes exhibiting global misalignment. Fig.3(c) shows an example of DLI and NLI aligned-pairs.
@@
LSD Composition. Following expert evaluation, we selected 6,425 aligned and in-focus scenes from the collected data to construct the LSD dataset, comprising high-resolution image pairs ranging from $3820\times 2160$ to $4080\times 3060$. The final dataset includes 2,893 indoor and 3,532 outdoor scenes captured under a broad range of illumination conditions. We categorize lighting into extreme low-light (0.1–50 lux) and low-light (50–200 lux) regimes. A total of 6,025 pairs are used for training, while 400 complete scenes are reserved for benchmarking. These benchmark scenes include both DLI and NLI variants across diverse lighting conditions and scene types, as illustrated in Fig. 3(b). Further dataset details are provided in the supplementary material.
@@
Figure 4: Distribution comparison between filtered and raw training images. (a) sharpness improvement. (b) intensity filtering
@@ 2.2 Refining Training Pairs
Despite controlled references, minor spatial degradations like defocus, local darkness, or overexposed highlights may persist [17, 22]. These issues can occasionally cause deep models to produce blurry regions and inconsistent exposure in extreme cases (see Sec. 4.4). To improve training reliability, we introduce a reference-guided patch pair elimination method. Thus, we extracted $512\times 512$ patches, yielding 208,853 pairs, then filter-out  14% using two filters:
(1) Brightness threshold. We discarded patches whose reference images had average intensity below $T=10$, computed as
$I_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}I_{i}$
This threshold filters out extremely dark reference patches that appear nearly black on standard displays and typically lack meaningful structure due to underexposure or ISP-induced smoothing. Removing these patches ensures the model learns from visually informative and reliable supervision.
(2) Confidence filter. We developed a VGG-based [23] scene classifier to evaluate patch quality evaluation. The confidence score $C$ for each reference image, computed as $C=f(\mathbf{I_{G}})$ (where $\mathbf{I_{G}}$ represents the reference image), was used to filter out pairs with $C<0.90$. Since reference images guide deep models in generating plausible outputs, we prioritized their quality during the filtering process. Fig. 4 compares the distributions of the raw and filtered patches, highlighting the improvements in intensity and sharpness in the refined training set.
@@ 2.3 LSD-U: Unpair Benchmarking Dataset
SLLIE is fundamentally an aesthetic task, requiring methods to generalize across diverse real-world scenarios. Accordingly, evaluation should be performed on datasets with broad scene variety and data distributions, rather than those limited by specific biases. Thus, we introduce LSD-U, including 2,117 unique unpaired test cases. Our benchmark features data from DSLR cameras, smartphones, low-light video frames, and social media images sourced from various users with different photographic styles and setups compared to the proposed LSD. The LSD-U distribution introduces varied biases, enabling SLLIE methods to be tested across a wide range of real-world conditions, including enhancing trillions of social media images.
@@ 3 Learning with TFFormer
TFFormer learns to map $\mathrm{T}\!:\mathbf{I}_{LL}\!\rightarrow\!(\mathbf{I}_{rec},\mathbf{I}_{ref})$, where $\mathbf{I}_{LL}$ is the input low-light image, $\mathbf{I}_{rec}$ an intermediate reconstruction, and $\mathbf{I}_{ref}$ the final refined output. The pipeline includes: (1) LC mapping for LC decomposition, (2) LC encoders with attention-based encoding, (3) cross-attention-based LC feature fusion, (4) a joint decoder generating $\mathbf{I}_{rec}$, and (5) an LC Guided Reconstruction Block (LCGRB) refining $\mathbf{I}_{rec}$ into $\mathbf{I}_{ref}$. This modular design facilitates structured fusion and progressive refinement. Figure 5 illustrates the pipeline.